# Zens, Díaz, Durante and Patacchini

### v1

@@ Abstract
Motivated by the above considerations, we rely on a separable construction that assumes $\bm{\Theta}={\bf B}\otimes{\bf A}$ and $\bm{\Sigma}=\bm{\Sigma_{B}}\otimes\bm{\Sigma_{A}}$, where $\otimes$ denotes the Kronecker product. In this factorization, ${\bf A}$ and ${\bf B}$ are $K\times K$ and $Q\times Q$ matrices that capture the average inter-temporal dependence between the $K$ crime categories within spatial locations, and among these $Q$ spatial locations within each crime category, respectively. Similarly, $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$ correspond to the $K\times K$ and $Q\times Q$ covariance matrices encoding the across-categories and across-locations instantaneous dependencies in reaction to shocks, respectively. In such a way, it is possible to disentangle spatial and crime-category dependencies, thereby enhancing interpretability, while preserving flexibility by letting such dependencies interact in the definition of $\bm{\Theta}$ and $\bm{\Sigma}$. Separable constructions have been successfully employed in black-box methods (e.g., Wang et al., 2025) to improve predictive accuracy, but the potential of these perspectives to allow for interpretable model-based inference on the complex dependencies within the criminal activity tensor have not yet been explored.
@@
${\bf Y}_{t}={\bf A}{\bf Y}_{t-1}{\bf B}^{\intercal}+{\bf E}_{t},\qquad{\bf E}_{t}\stackrel{{\scriptstyle\mbox{\footnotesize i.i.d.}}}{{\sim}}\mbox{MN}_{K,Q}({\bf 0},\bm{\Sigma}_{A},\bm{\Sigma}_{B}),$ (1)
for every $t=1,\ldots,T$. This formulation specializes the general class of multilinear tensor regression models in Hoff (2015) to our specific setting by considering a matrix-valued outcome related to a matrix-variate input via a temporal autoregressive structure; see also Billio et al. (2023) and the references therein for more general multi-way tensor autoregression constructions with a main focus on econometric applications. Crucially, compared to a fully-unstructured vector autoregressive formulation, the model in (1) reduces the number of parameters from $(KQ)^{2}+(KQ)(KQ+1)/2$ to $K^{2}+Q^{2}+K(K+1)/2+Q(Q+1)/2$. In our application to Chicago criminal activity data with $K=28$ crime categories and $Q=22$ spatial locations, this implies a reduction from $569{,}492$ parameters to $1{,}927$ (about three hundred times lower than those within the unstructured model). Interestingly, as illustrated in Section 5, such a remarkable reduction in complexity still yields a predictive performance competitive with routine-use black-box methods, meaning that model (1) properly accounts for a rich set of dynamics within and between crime categories, locations and times.
Although the above formulation achieves a sensible balance between predictive potential and interpretable inference on the dependencies in the criminal activity tensor, in practice, unrestricted estimation of the $K\times K$ and $Q\times Q$ matrices ${\bf A}$ and ${\bf B}$ may still result in a overly-parametrized model. In addition, as discussed in Section 2, it is reasonable to expect that both ${\bf A}$ and ${\bf B}$ would admit lower-dimensional representations due the sparse and structured inter-temporal dependencies associated with both crime categories and spatial locations. To account for this information and further reduce complexity, we combine model (1) with a low-rank factorization for ${\bf A}$ and ${\bf B}$ defined as
@@
Despite the potential of the above ideas for learning structure in three-way criminal activity tensors, to our knowledge formulations of this type have been overlooked within criminology settings. To cover such a gap we employ the model in (1)–(2) and perform inference on its parameters under a Bayesian perspective (see Section 4). Besides introducing additional shrinkage effects through prior distributions, this choice provides a fully probabilistic framework enabling formal uncertainty quantification, also in forecasting, and possible inclusion of expert information along with nuanced hierarchical structures for more sophisticated crime modeling exercises.
@@
To perform Bayesian inference under model (1)–(2) it is necessary to elicit prior distributions for its parameters, namely $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, ${\bf L}_{A}$, ${\bf L}_{B}$, ${\bf Z}_{A}$ and ${\bf Z}_{B}$. These priors are specified to facilitate tractable posterior inference, while introducing mild shrinkage effects that, however, do not overwhelm the information provided by the observed criminal activity tensor through the likelihood of model (1)–(2).
Consistent with the above goals, we follow routine-practice in Bayesian modeling of covariance matrices (e.g., Gelman et al., 2013) by relying on inverse-Wishart priors for $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$; namely
$\bm{\Sigma}_{A}\sim\mbox{inv-Wishart}(4,2\cdot{\bf I}_{K}),\qquad\quad\bm{\Sigma}_{B}\sim\mbox{inv-Wishart}(4,2\cdot{\bf I}_{Q}).$
Combined with the Gaussian likelihood induced by model (1)–(2), these priors admit conditional-conjugacy properties that facilitate Bayesian inference on the instantaneous dependence structures among the $K$ crime categories and between the $Q$ spatial locations. Notice that, although the hyperparameters of the above priors can be possibly set at different values, in the application in Section 5 we found these choices effective both in inference and in out-of-sample prediction.
As for the priors on ${\bf Z}_{A}$ and ${\bf Z}_{B}$, we adapt to our specific low-rank bilinear autoregressive setting the elicitation suggested for the general class of multilinear tensor regression models in Hoff (2015). To this end, we let
@@
thereby inducing via the low-rank construction in (2) weakly-diffuse and conditionally-conjugate priors on the separable inter-temporal effects in ${\bf A}$ and ${\bf B}$. Under (2), such effects are also controlled by ${\bf L}_{A}$ and ${\bf L}_{B}$ for which we assume flat priors combined with a processing step within the proposed Gibbs sampler that enforces the constraints ${\bf L}_{A}^{\intercal}{\bf L}_{A}={\bf I}_{R_{A}}$ and ${\bf L}_{B}^{\intercal}{\bf L}_{B}={\bf I}_{R_{B}}$, respectively. This choice implicitly fixes the scale of the elements within ${\bf L}_{A}$ and ${\bf L}_{B}$ thereby addressing rotational and scaling identification issues that may arise from the factorization in (2). While these identifiability problems only affect ${\bf L}_{A}$, ${\bf Z}_{A}$ and ${\bf L}_{B}$, ${\bf Z}_{B}$, and not the main objects of inference, i.e., ${\bf A}$ and ${\bf B}$, considering processing steps that address these potential issues improves mixing and convergence of the Gibbs sampler we design in Section 4.2 for posterior inference.
4.2 4.2 Posterior computation via Gibbs sampler
The prior distributions in Section 4.1 combined with the likelihood of model (1)–(2) yield a posterior for the parameters $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, ${\bf L}_{A}$, ${\bf L}_{B}$, ${\bf Z}_{A}$ and ${\bf Z}_{B}$ which is not available in closed form. Nonetheless, as clarified below, tractable full conditionals for each of these parameters given the others and the observed data are available, thereby allowing for Monte Carlo inference leveraging samples from the posterior of these parameters obtained via a Gibbs sampling algorithm.
@@
$\text{vec}(\tilde{{\bf Y}}_{(A)})=({\tilde{\bf X}}_{(A)}^{\intercal}{\bf Z}_{A}^{\intercal}\otimes{\bf I}_{K})\text{vec}({\bf L}_{A})+\text{vec}(\tilde{{\bf E}}_{(A)}),\quad\text{vec}(\tilde{{\bf E}}_{(A)})\sim\mbox{N}_{KTQ}({\bf 0},{\bf I}_{TQ}\otimes\bm{\Sigma}_{A}),$ (5)
thereby recasting model (1)–(2) as a Gaussian linear regression with coefficients $\text{vec}({\bf Z}_{A})$ (see (4)) or $\text{vec}({\bf L}_{A})$ (see (5)), respectively, conditionally on the other parameters. Combined with the priors for $\text{vec}({\bf Z}_{A})$ and $\text{vec}({\bf L}_{A})$ presented in Section 4.1, this formulation yields Gaussian full conditionals for both vectors. More specifically, $(\text{vec}({\bf Z}_{A})\mid-)\sim\mbox{N}_{R_{A}K}(\bm{\mu}_{{\bf Z}_{A}},\bm{\Sigma}_{{\bf Z}_{A}})$ with
@@
Leveraging the above samples for $\text{vec}({\bf Z}_{A})$ and $\text{vec}({\bf L}_{A})$ it is possible to obtain directly a draw from the full-conditional distribution of the matrix ${\bf A}$ by computing ${\bf A}={\bf L}_{A}{\bf Z}_{A}$ at these samples. Conditioned on ${\bf A}$, it is then possible to leverage the conditional conjugacy of the inverse-Wishart prior for $\bm{\Sigma}_{A}$ under model (3), to obtain
$(\bm{\Sigma}_{A}\mid-)\sim\mbox{inv-Wishart}(4+TQ,2\cdot{\bf I}_{K}+(\tilde{{\bf Y}}_{(A)}-{\bf A}{\tilde{\bf X}}_{(A)})(\tilde{{\bf Y}}_{(A)}-{\bf A}{\tilde{\bf X}}_{(A)})^{\intercal}).$ (8)
A similar reasoning and derivations can be also considered to obtain closed-form and tractable full conditionals for ${\bf Z}_{B}$, ${\bf L}_{B}$ and $\bm{\Sigma}_{B}$. More specifically, given ${\bf A}$ and $\bm{\Sigma}_{A}$, it is possible to obtain the following $B$-matricization of (1)
parameterized by $\text{vec}({\bf L}_{B})$. Combining (10) with the conditionally-conjugate Gaussian prior for ${\bf Z}_{B}$, implies that $(\text{vec}({\bf Z}_{B})\mid-)\sim\mbox{N}_{R_{B}Q}(\bm{\mu}_{{\bf Z}_{B}},\bm{\Sigma}_{{\bf Z}_{B}})$ with
$(\bm{\Sigma}_{B}\mid-)\sim\mbox{inv-Wishart}(4+TK,2\cdot{\bf I}_{Q}+(\tilde{{\bf Y}}_{(B)}-{\bf B}{\tilde{\bf X}}_{(B)})(\tilde{{\bf Y}}_{(B)}-{\bf B}{\tilde{\bf X}}_{(B)})^{\intercal}).$ (14)
@@
Algorithm 1 summarizes the steps of the Gibbs sampler resulting from the above derivations. Notice that, due to the Kronecker structure of the matrix autoregressive model and the low-rank factorization of the coefficient matrices, some identifiability issues may arise. First, the matrix products $\mathbf{L}_{A}\mathbf{Z}_{A}$ and $\mathbf{L}_{B}\mathbf{Z}_{B}$ are only identifiable up to rotation, sign, scale and ordering of the respective rows and columns. Enforcing the constraints ${\bf L}_{A}^{\intercal}{\bf L}_{A}={\bf I}_{R_{A}}$ and ${\bf L}_{B}^{\intercal}{\bf L}_{B}={\bf I}_{R_{B}}$ discussed in Section 4.1 ensures scale and rotational identification, and increases numerical stability of the algorithm. Such constraints are included within the Gibbs sampler via an intermediate processing step that first computes the QR decomposition of both ${\bf L}_{A}$ and ${\bf L}_{B}$, i.e., ${\bf L}_{A}={\bf Q}_{L_{A}}{\bf R}_{L_{A}}$ and ${\bf L}_{B}={\bf Q}_{L_{B}}{\bf R}_{L_{B}}$, and then redefines ${\bf L}_{A}={\bf Q}_{L_{A}}$, ${\bf Z}_{A}={\bf R}_{L_{A}}{\bf Z}_{A}$, and ${\bf L}_{B}={\bf Q}_{L_{B}}$, ${\bf Z}_{B}={\bf R}_{L_{B}}{\bf Z}_{B}$. In such a way, ${\bf A}={\bf L}_{A}{\bf Z}_{A}$ and ${\bf B}={\bf L}_{B}{\bf Z}_{B}$ remain unchanged, while meeting the identifiability constraints ${\bf L}_{A}^{\intercal}{\bf L}_{A}={\bf I}_{R_{A}}$ and ${\bf L}_{B}^{\intercal}{\bf L}_{B}={\bf I}_{R_{B}}$. Column switching and sign switching was not an issue for the considered exercises, but could be readily resolved using order constraints during the mcmc or via ex post identification strategies. Finally, recalling Hoff (2015), ${\bf A}$, ${\bf B}$ and ${\bf\Sigma}_{A}$, ${\bf\Sigma}_{B}$ are only identified up to scaling, i.e., ${\bf B}\otimes{\bf A}=({\bf B}/c)\otimes(c{\bf A})$ and ${\bf\Sigma}_{B}\otimes{\bf\Sigma}_{A}=({\bf\Sigma}_{B}/c)\otimes(c{\bf\Sigma}_{A})$. To address this issue, we follow the recommendation within Appendix B of Hoff (2015) by rescaling ${\bf A}$, ${\bf B}$ and ${\bf\Sigma}_{A}$, ${\bf\Sigma}_{B}$ to maintain a constant relative magnitude among $||{\bf A}||^{2}$ and $||{\bf B}||^{2}$, and between $||{\bf\Sigma}_{A}||^{2}$ and $||{\bf\Sigma}_{B}||^{2}$, respectively.
@@
We conclude by illustrating the predictive performance and the inference potential of the Bayesian low-rank bilinear autoregressive model in Sections 34, with a focus on criminal activity data in the city of Chicago. These data are described in detail in Section 5.1. The results of the analysis under the proposed model can be found in Section 5.2.
@@
The three-way criminal activity tensor we rely on is available at https://www.chicagopolice.org/data-statistics/ and comprises time-stamped reports from the Chicago Police Department on $K=28$ crime categories within the $Q=22$ districts of the city. In our application, these data are aggregated on a monthly timescale from January 2001 until December 2019 in order to avoid structural breaks due to the onset of the covid-19 pandemic in the early 2020. Figure 1 provides a graphical representation of the dynamic log-counts of these reports for each crime category, aggregated over districts, whereas Figure 2 displays selected heatmaps of the number of crime reports across the $Q=22$ districts for specific categories, averaged over the time window analyzed. Consistent with the discussion in Section 2, both figures clarify that criminal activities are characterized by heterogeneous structures across time, crime categories and districts, which further showcase nuanced and complex interdependence patterns, thereby motivating the implementation of the model developed in Sections 34.
Notice that the $K=28$ crime categories we study in this application mainly follow the classification system employed by the Federal Bureau of Investigation (fbi), which aligns with the National Incident-Based Reporting System (nibrs). However, we observed that a nibrs category Miscellaneous Other often played an intriguing role in our results. To enhance the interpretability and specificity of our analysis, we therefore decided to redistribute crimes within this category based on the corresponding Illinois Uniform Crime Reporting (iucr) code. For example, we reallocated reports corresponding to iucr code 2092 – solicit narcotics on public way to the fbi category Drug Abuse (18), which therefore represents a broader, narcotics-related category. Furthermore, certain other iucr codes previously part of fbi category Miscellaneous Other have been bundled into new, separate, crime categories. An example for this are several codes related to threat and harassment, for instance 2820 - telephone threat or 2826 - harassment by electronic means. We reclassified these codes into an explicit Threat & Harassment category. This refinement of the crime classification allows us to capture the nuances of different crime types more accurately and facilitates a more in depth analysis.
@@
Figure 2: Heatmaps of reports’ counts for selected crime categories in each district of Chicago, averaged over time. Data are shown on a logarithmic scale.
Before discussing the results of our analysis, it is also worth to mention that any study based on police reports provides only a partial understanding of the underlying criminal activity patterns due to several factors. First, police reports inherently represent the information on criminal activity filtered through the lens of law enforcement. As such, it is important to acknowledge that the reported incidents may not perfectly align with the actual criminal activity across the city. Second, the spatial dimension of these reports introduces complexities, as different neighborhoods may have varying levels of trust in the police, leading to differences in reporting patterns. Finally, the type of crime being reported and the temporal aspect can also impact the results, further highlighting the need for caution when interpreting the findings. Nonetheless, despite these limitations, analyzing police reports remains the standard approach in state-of-the-art literature, and is commonly used for forecasting and utilized by the law enforcement itself. It provides valuable insights into crime patterns, police activity, and the structures that emerge from the interaction among the two. That is, the results may be influenced by both the system of policing and the actual criminal activity, making it challenging to disentangle the two. Nonetheless, the interpretable model-based perspective we propose can uncover hidden patterns within the data produced by law enforcement, thereby providing police with more refined methods to examine and leverage information from these reports.

### v2

@@ Abstract
Motivated by the above considerations, we rely on a separable construction that assumes $\bm{\Theta}={\bf B}\otimes{\bf A}$ and $\bm{\Sigma}=\bm{\Sigma_{B}}\otimes\bm{\Sigma_{A}}$, where $\otimes$ denotes the Kronecker product. In this factorization, $\bf A$} and $\bf B$} are $K\times K$ and $Q\times Q$ matrices that capture the “average” inter-temporal dependence between the $K$ crime categories within spatial locations, and among these $Q$ spatial locations within each crime category, respectively. Similarly, $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$ correspond to the $K\times K$ and $Q\times Q$ covariance matrices encoding the across-categories and across-locations instantaneous dependencies in reaction to shocks, respectively. In such a way, it is possible to disentangle spatial and crime-category dependencies, thereby enhancing interpretability, while preserving flexibility by letting such dependencies interact in the definition of $\bm{\Theta}$ and $\bm{\Sigma}$. Separable constructions have been successfully employed in black-box methods (e.g., Wang et al., 2025) to improve predictive accuracy, but the potential of these perspectives to allow for interpretable model-based inference on the complex dependencies within the criminal activity tensor have not yet been explored.
@@
${\bf Y}_{t}={\bf A}{\bf Y}_{t-1}{\bf B}^{\intercal}+{\bf E}_{t},\qquad{\bf E}_{t}\sim\mbox{MN}_{K,Q}({\bf 0},\bm{\Sigma}_{A},\bm{\Sigma}_{B}),$ (1)
independently for each $t=1,\ldots,T$, where ${\bf E}_{t}$ is the $K\times Q$ matrix with noise entries $\varepsilon_{kqt}$, while ${\bf E}_{t}\sim\mbox{MN}_{K,Q}({\bf 0},\bm{\Sigma}_{A},\bm{\Sigma}_{B})$ implies $\mbox{vec}({\bf E}_{t})=\bm{\varepsilon}_{t}\sim\mbox{N}_{KQ}({\bf 0},\bm{\Sigma_{B}}\otimes\bm{\Sigma_{A}})$. Therefore, under (1), the model for each entry $y_{kqt}$ of ${\bf Y}_{t}$ is $y_{kqt}=\sum_{k^{\prime}=1}^{K}\sum_{q^{\prime}=1}^{Q}a_{kk^{\prime}}b_{qq^{\prime}}y_{k^{\prime}q^{\prime}(t-1)}+\varepsilon_{kqt}$, where $a_{kk^{\prime}}$ and $b_{qq^{\prime}}$ denote the entries of position $(k,k^{\prime})$ and $(q,q^{\prime})$ of $\bf A$} and $\bf B$}, respectively. This formulation specializes the general class of multilinear tensor regression models proposed by Hoff (2015) [^1] Notice that while the original focus of Hoff (2015) is on dynamic networks where ${\bf Y}_{t}$ denotes an adjacency matrix encoding connectivity data, the formulation proposed extends to general tensor-valued data, beyond dynamic networks. In our setting ${\bf Y}_{t}$ does not measure directed relationships, but rather crime levels at time $t$, for each combination of crime category $k$ and spatial location $q$, denoting, respectively, the rows and columns of ${\bf Y}_{t}$. As such, the ordering within rows and columns does not play a role in the model formulation and results. to our specific setting by considering a matrix-valued outcome related to a matrix-variate input through a temporal autoregressive structure; see also Martinez-Beneito et al. (2017), Billio et al. (2023) and the references therein for general multi-way tensor regression constructions with a focus on disease mapping and econometric applications, respectively. Crucially, compared to a fully-unstructured vector autoregressive formulation, the statistical model in (1) reduces the number of parameters from $(KQ)^{2}+(KQ)(KQ+1)/2$ to $K^{2}+Q^{2}+K(K+1)/2+Q(Q+1)/2$. In our application to Chicago criminal activity data with $K=28$ crime categories and $Q=22$ spatial locations, this implies a reduction from 569,492569{,}492 parameters to 1,9271{,}927 (about three hundred times lower than those within the unstructured model). Interestingly, as illustrated in Section 5, such a remarkable reduction in complexity still yields a predictive performance competitive with routine-use black-box methods, meaning that the model in (1) properly accounts for a rich set of dynamics within and between crime categories, locations and times.
Although the above formulation achieves a sensible balance between predictive potential and interpretable inference on the dependencies in the criminal activity tensor, in practice, unrestricted estimation of the $K\times K$ and $Q\times Q$ matrices $\bf A$} and $\bf B$} may still result in a overly-parametrized model. In addition, as discussed in Section 2, it is reasonable to expect that both $\bf A$} and $\bf B$} admit lower-dimensional representations due to the sparse and structured inter-temporal dependencies associated with both crime categories and spatial locations. To account for this information and further reduce complexity, we combine (1) with a low-rank factorization for $\bf A$} and $\bf B$} defined as
@@
Despite the potential of the above ideas for learning structure within three-way criminal activity tensors, to our knowledge formulations of this type have been underutilized in criminology. To cover such a gap we employ the model in (1)–(2) and perform inference on its parameters under a Bayesian perspective (see Section 4). This choice provides a fully probabilistic framework enabling formal uncertainty quantification, also in forecasting, and possible inclusion of expert information along with nuanced hierarchical structures for more sophisticated crime modeling exercises.
@@
To perform Bayesian inference under model (1)–(2) it is necessary to elicit prior distributions for its parameters, namely $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, ${\bf L}_{A}$, ${\bf L}_{B}$, ${\bf Z}_{A}$ and ${\bf Z}_{B}$. These priors are specified to facilitate tractable posterior inference, while favoring weakly-informative or non-informative elicitations which do not overwhelm the information provided by the observed criminal activity tensor through the likelihood of model (1)–(2).
Consistent with the above goals, we follow routine practice in Bayesian modeling of covariance matrices (e.g., Gelman et al., 2013) by relying on inverse-Wishart priors for $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$; namely
$\bm{\Sigma}_{A}\sim\mbox{inv-Wishart}(2,2\cdot{\bf I}_{K}),\qquad\quad\bm{\Sigma}_{B}\sim\mbox{inv-Wishart}(2,2\cdot{\bf I}_{Q}),$
where the hyperparameters are selected to induce a diffuse specification that proved effective both in inference and in out-of-sample prediction for the application in Section 5 (when tested in sensitivity analyses, moderate changes of these hyperparameters did not substantially affect the final results). Combined with the Gaussian likelihood induced by the model in (1)–(2), the above priors admit conditional-conjugacy properties that facilitate Bayesian inference on the instantaneous dependence structures among the $K$ crime categories and between the $Q$ spatial locations. Notice that these priors do not include specific structures within $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$, but rather learn flexibly both matrices from the data. This choice might appear suboptimal for $\bm{\Sigma}_{B}$, which measures dependence across locations, thereby suggesting alternative representations that explicitly include spatial proximity information regulating these dependencies, such as car–type constructions (e.g., Banerjee et al., 2015). Although it would be interesting to extend the proposed framework to explicitly incorporate such information, this perspective would undermine the flexibility of the model in our settings. In fact, as anticipated in Section 2, the spatial dependencies in criminal activities typically arise from the combined effect of several underlying factors such as localized presence of criminal organizations or distinctive policing strategies, which are often difficult to include in the model only via spatial proximity information. Avoiding such constraints allows us to learn in Section 5 (see Figures 4 and 5) dependencies among Chicago districts that partially depart from pure spatial proximity and provide quantitative evidence supporting the reported presence of organized crime and structured gangs activities in the city (e.g., Aspholm, 2019; Schnell et al., 2017).
As for the priors on ${\bf Z}_{A}$ and ${\bf Z}_{B}$, we adapt to our specific low-rank bilinear autoregressive setting the elicitation suggested for the general class of multilinear tensor regression models in Hoff (2015). In particular, we assume independent $\mbox{N}(0,10)$ priors for all the entries of ${\bf Z}_{A}$ and ${\bf Z}_{B}$, where the variance of $10$ yields a weakly-diffuse specification that introduces some regularization effect, but avoids imposing excessive information through the prior. In matrix form, this choice implies
@@
thereby inducing via the low-rank construction in (2) weakly-diffuse and conditionally-conjugate priors also on the separable inter-temporal effects in $\bf A$} and $\bf B$}. Under (2), such effects are also controlled by ${\bf L}_{A}$ and ${\bf L}_{B}$. Consistent with the previous elicitations, the entries of such matrices are assigned non-informative flat priors (i.e., uniform in $\mathbb{R}$) further combined with a processing step within the proposed Gibbs sampler that enforces the constraints ${\bf L}_{A}^{\intercal}{\bf L}_{A}={\bf I}_{R_{A}}$ and ${\bf L}_{B}^{\intercal}{\bf L}_{B}={\bf I}_{R_{B}}$, respectively. This choice implicitly fixes the scale of the elements within ${\bf L}_{A}$ and ${\bf L}_{B}$, thereby addressing rotational and scaling identification issues that may arise from the factorization in (2). While these identifiability problems only affect ${\bf L}_{A}$, ${\bf Z}_{A}$ and ${\bf L}_{B}$, ${\bf Z}_{B}$, and not the main objects of inference, i.e., $\bf A$} and $\bf B$}, considering processing steps that address these potential issues improves mixing and convergence of the Gibbs sampler we design in Section 4.2 for posterior inference.
4.2 4.2 Posterior computation via Gibbs sampling
The prior distributions in Section 4.1 combined with the likelihood of model (1)–(2) yield a posterior for the parameters $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, ${\bf L}_{A}$, ${\bf L}_{B}$, ${\bf Z}_{A}$ and ${\bf Z}_{B}$ which is not available in closed form. Nonetheless, tractable full conditionals for each of these parameters, given the others and the observed data are available, thereby allowing for Monte Carlo inference leveraging samples from the posterior of these parameters obtained via a Gibbs sampling algorithm.
@@
$\text{vec}(\tilde{{\bf Y}}_{(A)})=({\tilde{\bf X}}_{(A)}^{\intercal}{\bf Z}_{A}^{\intercal}\otimes{\bf I}_{K})\text{vec}({\bf L}_{A})+\text{vec}(\tilde{{\bf E}}_{(A)}),\quad\text{vec}(\tilde{{\bf E}}_{(A)})\sim\mbox{N}_{KTQ}({\bf 0},{\bf I}_{TQ}\otimes\bm{\Sigma}_{A}).$ (5)
This allows to recast model (1)–(2) as a Gaussian linear regression with coefficients $\text{vec}({\bf Z}_{A})$ (see (4)) or $\text{vec}({\bf L}_{A})$ (see (5)), respectively, conditionally on the other parameters. Combined with the priors for $\text{vec}({\bf Z}_{A})$ and $\text{vec}({\bf L}_{A})$ presented in Section 4.1, this formulation yields Gaussian full conditionals for both vectors. More specifically, $(\text{vec}({\bf Z}_{A})\mid-)\sim\mbox{N}_{R_{A}K}(\bm{\mu}_{{\bf Z}_{A}},\bm{\Sigma}_{{\bf Z}_{A}})$ with
@@
Leveraging the above samples for $\text{vec}({\bf Z}_{A})$ and $\text{vec}({\bf L}_{A})$, a draw from the full conditional of $\bf A$} can be obtained by computing ${\bf A}={\bf L}_{A}{\bf Z}_{A}$ at these samples. Conditioned on $\bf A$}, it is then possible to leverage the conditional conjugacy of the inverse-Wishart prior for $\bm{\Sigma}_{A}$ under (3), to obtain
$(\bm{\Sigma}_{A}\mid-)\sim\mbox{inv-Wishart}(2+TQ,2\cdot{\bf I}_{K}+(\tilde{{\bf Y}}_{(A)}-{\bf A}{\tilde{\bf X}}_{(A)})(\tilde{{\bf Y}}_{(A)}-{\bf A}{\tilde{\bf X}}_{(A)})^{\intercal}).$ (8)
@@
A similar reasoning and derivations apply to the full conditionals for ${\bf Z}_{B}$, ${\bf L}_{B}$ and $\bm{\Sigma}_{B}$. More specifically, given $\bf A$} and $\bm{\Sigma}_{A}$, it is possible to obtain the following $B$-matricization of (1)
parameterized by $\text{vec}({\bf L}_{B})$.
Combining (10) with the conditionally conjugate matrix-variate normal prior for ${\bf Z}_{B}$, implies that $(\text{vec}({\bf Z}_{B})\mid-)\sim\mbox{N}_{R_{B}Q}(\bm{\mu}_{{\bf Z}_{B}},\bm{\Sigma}_{{\bf Z}_{B}})$ with
$(\bm{\Sigma}_{B}\mid-)\sim\mbox{inv-Wishart}(2+TK,2\cdot{\bf I}_{Q}+(\tilde{{\bf Y}}_{(B)}-{\bf B}{\tilde{\bf X}}_{(B)})(\tilde{{\bf Y}}_{(B)}-{\bf B}{\tilde{\bf X}}_{(B)})^{\intercal}).$ (14)
Algorithm 1 summarizes the steps of the Gibbs sampler resulting from the above derivations. Notice that, due to the Kronecker structure of the matrix autoregressive model and the low-rank factorization of the coefficient matrices, some identifiability issues may arise. First, the matrix products $\mathbf{L}_{A}\mathbf{Z}_{A}$ and $\mathbf{L}_{B}\mathbf{Z}_{B}$ are only identifiable up to rotation, sign, scale and ordering of the respective rows and columns. Enforcing the constraints ${\bf L}_{A}^{\intercal}{\bf L}_{A}={\bf I}_{R_{A}}$ and ${\bf L}_{B}^{\intercal}{\bf L}_{B}={\bf I}_{R_{B}}$ discussed in Section 4.1 ensures scale and rotational identification, and increases numerical stability of the algorithm. Such constraints are included within the Gibbs sampler via an intermediate processing step that first computes the QR decomposition of both ${\bf L}_{A}$ and ${\bf L}_{B}$, i.e., ${\bf L}_{A}={\bf Q}_{L_{A}}{\bf R}_{L_{A}}$ and ${\bf L}_{B}={\bf Q}_{L_{B}}{\bf R}_{L_{B}}$, and then redefines ${\bf L}_{A}={\bf Q}_{L_{A}}$, ${\bf Z}_{A}={\bf R}_{L_{A}}{\bf Z}_{A}$, and ${\bf L}_{B}={\bf Q}_{L_{B}}$, ${\bf Z}_{B}={\bf R}_{L_{B}}{\bf Z}_{B}$. In such a way, ${\bf A}={\bf L}_{A}{\bf Z}_{A}$ and ${\bf B}={\bf L}_{B}{\bf Z}_{B}$ remain unchanged, while meeting the identifiability constraints ${\bf L}_{A}^{\intercal}{\bf L}_{A}={\bf I}_{R_{A}}$ and ${\bf L}_{B}^{\intercal}{\bf L}_{B}={\bf I}_{R_{B}}$. Column and sign switching was not an issue in our case, but could be readily resolved via order constraints within the mcmc or ex post identification strategies. Finally, recalling Hoff (2015), $\bf A$}, $\bf B$} and ${\bf\Sigma}_{A}$, ${\bf\Sigma}_{B}$ are only identified up to scaling, i.e., ${\bf B}\otimes{\bf A}=({\bf B}/c)\otimes(c{\bf A})$ and ${\bf\Sigma}_{B}\otimes{\bf\Sigma}_{A}=({\bf\Sigma}_{B}/c)\otimes(c{\bf\Sigma}_{A})$. To address this issue, we follow the recommendation in Appendix B of Hoff (2015) by rescaling $\bf A$}, $\bf B$} and ${\bf\Sigma}_{A}$, ${\bf\Sigma}_{B}$ to maintain a constant relative magnitude among ${\bf A}||^{2}$ and ${\bf B}||^{2}$, and between ${\bf\Sigma}_{A}||^{2}$ and ${\bf\Sigma}_{B}||^{2}$, respectively.
@@
We proceed by illustrating the predictive performance and the inference potential of the model proposed in Sections 3 and 4, with a focus on criminal activity data in Chicago. These data are described in Section 5.1. The results of the analysis can be found in Section 5.2.
@@
The three-way criminal activity tensor we rely on is available at https://www.chicagopolice.org/data-statistics/ and comprises time-stamped reports from the Chicago Police Department on $K=28$ crime categories within the $Q=22$ districts of the city. In our application, these data are aggregated on a monthly timescale from January 2001 until December 2019 in order to avoid structural breaks due to the covid-19 pandemic in the early 2020. Figure 1 shows the dynamic log-counts of these reports for each crime category, aggregated over districts, while Figure 2 displays selected heatmaps of the log-transformed number of crime reports across the $Q=22$ districts for specific categories, averaged over the time window analyzed. Consistent with the discussion in Section 2, both figures clarify that criminal activities display heterogeneous structures across time, crime categories and districts, which further showcase complex interdependencies, thus motivating the implementation of the model developed in Sections 34. As discussed in Section 3, this model is applied to the de-trended and seasonally adjusted version of the time series in Figure 1.
Notice that the $K=28$ crime categories we study mainly follow the classification system employed by the Federal Bureau of Investigation (fbi), which aligns with the National Incident-Based Reporting System (nibrs). However, we observed that a nibrs category Miscellaneous Other often played an intriguing role in our results. To enhance interpretability, we therefore decided to redistribute crimes within this category based on the corresponding Illinois Uniform Crime Reporting (iucr) code (see the Supplementary Material). For example, we reallocated reports corresponding to iucr code 2092 – solicit narcotics on public way to the fbi category Drug Abuse (18), which therefore represents a broader, narcotics-related category. Furthermore, certain other iucr codes previously part of fbi category Miscellaneous Other have been bundled into new, separate, crime categories. An example for this are several codes related to threat and harassment, for instance 2820 - telephone threat or 2826 - harassment by electronic means. We reclassified these codes into an explicit Threat & Harassment category. This refinement of the crime classification allows us to capture the nuances of different crimes more accurately.
@@
Before discussing the results of our analysis, it is also worth mentioning that any study based on police reports provides only a partial understanding of the underlying criminal activity patterns due to several factors. First, police reports inherently represent the information on criminal activity filtered through the lens of law enforcement. As such, it is important to acknowledge that the reported incidents may not perfectly align with the actual criminal activity across the city. Second, the spatial dimension of these reports introduces complexities, as different neighborhoods may have varying levels of trust in the police, leading to differences in reporting patterns. Finally, the type of crime being reported and the temporal aspect can also impact the results, further highlighting the need for caution when interpreting the findings. Nonetheless, despite these limitations, analyzing police reports remains the standard approach in state-of-the-art literature, and is commonly used for forecasting and utilized by the law enforcement itself. It provides valuable insights into crime patterns, police activity, and the structures that emerge from the interaction among the two. That is, the results may be influenced by both the system of policing and the actual criminal activity, making it challenging to disentangle the two. Nonetheless, the interpretable model we propose can uncover hidden patterns in the data produced by law enforcement, thus providing police with more refined methods to leverage information from these reports.
Figure 2: Heatmaps of reports’ log-transformed counts for selected crime categories in each district of Chicago, averaged over time.